HUBBLE SPACE TELESCOPE DAILY REPORT #1552 PERIOD COVERED: 0000Z (UTC) 01/18/96 - 0000Z (UTC) 01/19/96 Daily Status Report as of 019/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/2 5869 (Chromospheric Activity in Quiet stars: Halo = Subdwarfs) The High Resolution Spectrograph (Side Two) was to be used to make = spectrographic observations of the quiet star HD106516 to study its = chromospheric activity. This is a search for Mg II emission in low = metallicity solar temperature field stars in the galactic halo. The guide = star acquisition for this proposal at 018/022648Z failed, an event = recorded in HSTAR 5467. No further problems were reported. 1.2 Completed WF/PC-2 5972 (Determination of the Extragalactic Distance = Scale) The WF/PC-2 was used to make observations of the galaxy NGC-2090 = looking for standard candle Cepheid variable stars. The purpose of these = observations is to help estimate the Hubble constant with an accuracy of = 10%. The observations were executed as scheduled, and no anomalies were = noted. 1.3 Completed WF/PC-2 6187 (WF/PC-2 Cycle 5 Calibration: Earth Flat) The WF/PC-2 was used to make Earth streak flat calibration images. = These streak flats will be used to produce flat fields, to map the OTA = illumination pattern and, in combination with other data, to produce new = superflats. The observations were completed as planned, and no problems = were reported. 1.4 Completed WF/PC-2 6136 (Positioning and Proper Motion Measurement = of PSR0656+14) The WF/PC-2 was used to make observations of radio pulsar = STAR-06594 (aka, PSR0656+14) to search for an optical counterpart and to = study its positioning and proper motion. The results will be compared to = existing ground based data. The observations were executed as scheduled, = and no anomalies were noted. 1.5 Completed Two Sets of FOS/RD 6236 (FOS Cycle 5: Location of Spectra = and Wavelength Calibration: Red Side Detector) The Faint Object Spectrograph (Red Side) was used to make = calibration observations in order to determine the location of the spectra = and to perform a wavelength calibration. All observations were internal. = A Taled calibration was also completed. The observations were completed = as planned, and no anomalies were reported. 1.6 Completed WF/PC-2 5949 (High Resolution Imaging of Galaxies = Responsible for QSO Lyman Alpha Absorption.) The WF/PC-2 was used to make high resolution observations of = Field-0405 in order to study galaxies which are responsible for = intermediate redshift QSO Lyman Alpha absorption. It is the extended = gaseous region around these galaxies that produces the absorption. The = observations were executed as scheduled, and no anomalies were noted. 1.7 Completed WF/PC-2 5983 (Curcumstellar Disks in Young Nebulous = Clusters) The WF/PC-2 was to be used to make observations of the young nearby = nebulous cluster NGC-2024-2 looking for possible curcumstellar disks = (Solar nebula analogs), like the type seen in the Orion Nebula. At = 018/131800Z, the guide star acquisition for this proposal failed. This = occurrence was reported in HSTAR 5469. No further problems were reported. 1.8 Completed WF/PC-2 6345 (Determining the Binary Frequency for = Ultra-Cool, Nearby M Dwarf Stars) The WF/PC-2 was used to make observations of an ultra-cool, nearby, = M class dwarf star LHS2243 to determine the frequency for it being binary = system and try to detect a stellar and sub-stellar companion. The = observations were executed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2 5505 (Deep Probes of Gravitational Lens Systems) The WF/PC-2 was used to make deep images of of the elliptical = galaxy MG1654+134 which is a member of the Coma Cluster. The intent is to = improve our understanding of the stellar populations in ellipticals and to = determine the distance to the Coma Cluster. About 16 minutes after a = normal guide star acquisition, there was an approximate one minute loss of = lock at 018/181730Z. This is documented in HSTAR 5470. Otherwise, this = 12-hour observation was completed as planned, and no further problems were = noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 5 As documented in HSTAR 5469, there was a failure in the FGS = acquisition at 018/131800. At 018/1730Z, there was an approximate one minute loss of lock = after a nominal acquisition sequence. Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 18 Successful : 18 2.3 Operations Notes: The battery #5 capacity test was begun at 018/1638Z. Thus far, = everything is proceeding niminally. On two occasions the command line from SAMS, the backup facility, = was utilized. The capability did not work on the first attempt, but = performed flawlessly on the second attempt. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH